OpenStudy (anonymous):

find the equation of a line passing though (2,-3)(6,5) and put it in ax=by=c. Help please and thank you (:

OpenStudy (anonymous):

first u need to find the slope delta y / delta x ... or y2-y1/x2-x1

OpenStudy (anonymous):

2 ?

OpenStudy (anonymous):

\[\frac{ 5-(-3) }{ 6-2 }=\frac{ 8 }{ 4 }=2=m\]

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

now plug in one coordinate and your slope of 2 into the point slope formula. y-y1 = m(x-x1)

OpenStudy (anonymous):

I am not sure what ax=by=c form is.

OpenStudy (anonymous):

i got it ! y-2x=-7. thank you ! (: appreciate it

OpenStudy (anonymous):

no problem :) good luck